Sue has a sock drawer where she collects nickels, dimes, and quarters. The total number of coins in her drawer is 37 with a total of $4.00. The number of dimes is 7 more than the number of nickels. How many coins of each type does she have?
Answer by TimothyLamb(4379)

---
x + y + z = 37
5x + 10y + 25z = 400
y = x + 7
---
put the system of linear equations into standard form
---
x + y + z = 37
5x + 10y + 25z = 400
x - y = -7
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= nickels = 12
y= dimes = 19
z= quarters = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php